# On the inadequacy of nudging data assimilation algorithms for non-dissipative systems: A Computational Examination of the Korteweg de-Vries and Lorenz equations

### v1

@@ 2. Preliminaries and Analytical Results
Moreover, because of the $\frac{L}{k}$ periodicity of $f(x)$ one has that $v(x,t)$ is also a solution of Equation 2.11. As $v(x,0)=u^{in}(x)$, then by the uniqueness of solutions to Equation 2.11 we obtain that $v(x,t)=u(x,t)$, hence we have that $u(x+\frac{L}{k},t)=u(x,t)$. The fact that $u$ restricted to $[-T,T]$ is continuous and bounded follows immediately from Theorem 10 in [13]. ∎
Corollary 2.0.4.
Under the conditions of 2.0.3 the space of periodic functions with period $\frac{L}{k}$ is invariant under the solution of Equation 2.11.
Proposition 2.0.5.
Suppose $k\in\mathbf{Z}\setminus\{0\}$ is given and suppose $u^{in}_{j}\in H^{2}_{\text{per}}$, for $j=1,2$, satisfying
$u^{in}_{j}(x+\frac{L}{k})=u_{j}^{in}(x)$ (2.13)
for a.e. $x\in\mathbf{R}$ and $u_{j}^{in}$ has spatial mean zero. Moreover, assume $\norm{u^{in}_{1}}_{L^{2}}\neq\norm{u^{in}_{2}}_{L^{2}}$ and let $u_{j}(x,t)$ be the corresponding solutions of Equation 2.6. Then
@@
    $u_{j}(x+\frac{L}{k},t)=u_{j}(x,t)$ for a.e. $x\in\mathbf{R}$ and for $j=1,2$ and all $t\in\mathbf{R}$.
@@
    $P_{M}u_{1}(\cdot,t)=P_{M}u_{2}(\cdot,t)$ for any positive integer $M<\absolutevalue{k},$ and every $t\in\mathbf{R}$.
@@
    $\limsup\limits_{t\to\infty}\norm{u_{1}(\cdot,t)-u_{2}(\cdot,t)}_{L^{2}}>0.$
@@
We note that as Equation 2.11 is a particular case of Equation 2.11 (i.e. $\gamma=0,\beta=0$), then (i) follows immediately from 2.0.3. Part (ii) then follows immediately from part (i) combined with 2.0.2.
Let us now prove part (iii). Let $u_{1}$ and $u_{2}$ be two solutions to Equation 2.6 under the assumptions of the proposition. Without loss of generality, suppose $\norm{u_{1}^{\text{in}}}_{L^{2}}>\norm{u_{2}^{\text{in}}}_{L^{2}}$. We note that the $L^{2}$ norm is a conserved quantity for Equation 2.6 and thus for $j=1,2$ we have that $\norm{u_{j}(t)}_{L^{2}}=\norm{u_{j}^{\text{in}}}_{L^{2}}$ holds for all $t\in\mathbf{R}$. The result follows from an application of the reverse triangle inequality as
$0<\norm{u_{1}(\cdot,t)}_{L^{2}}-\norm{u_{2}(\cdot,t)}_{L^{2}}\leq\norm{u_{1}(\cdot,t)-u_{2}(\cdot,t)}_{L^{2}}$ (2.14)
holds for all time $t\in\mathbf{R}$.
Remark 2.0.6.
2.0.5 can be adjusted without change to allow $u_{1}$ and $u_{2}$ to be initialized with periods $\frac{L}{k_{1}}$ and $\frac{L}{k_{2}}$, respectively. The only changes in this case are that $u_{j}$ remain $\frac{2}{k_{j}}$ periodic for all $t\in\mathbf{R}$ in (i) and $M<\min\{\absolutevalue{k_{1}},\absolutevalue{k_{2}}\}$ | } is required for (ii).
Now, using these propositions, we show that Equation 2.6 does not have the finitely many determining modes property.
Theorem 2.0.7.
The Korteweg de-Vries equation Equation 2.6 subject to periodic boundary conditions with period $L$ does not possess the finitely many determining modes property in $L^{2}$.
@@
The proof of this follows straightforwardly from 2.0.5. Assume, by way of contradiction, that the KdV equation enjoys the finitely many determining Fourier modes property in $L^{2}_{\text{per}}$. That is, there exists some $M\in\mathbf{N}$ such that for any two solutions $u_{1}$ and $u_{2}$ to Equation 2.6 that if
$\lim_{t\to\infty}\norm{P_{M}(u_{1}(t)-u_{2}(t))}_{L^{2}}=0,$ (2.15)
@@
$\lim_{t\to\infty}\norm{u_{1}(t)-u_{2}(t)}_{L^{2}}=0.$ (2.16)
Now, we will simply use $k=M+1$, and consider $u_{1}$ and $u_{2}$ to be any two solutions to Equation 2.6 such that $u_{1}$ and $u_{2}$ both are smooth enough and have zero spatial mean, are $\frac{L}{k}$ periodic initially, and $\norm{u_{1}^{in}}_{L^{2}}\neq\norm{u_{2}^{in}}_{L^{2}}$. We note that two such solutions must exist, as one could take $u_{1}$ to be the zero solution and $u_{2}$ to be the solution generated from the initial data $u_{2}^{in}=\cos(\frac{2\pi}{L}kx).$
Now, we immediately obtain a contradiction using 2.0.5. ∎
Remark 2.0.8.
The proof above can be adapted almost without change to the unforced 2D incompressible Euler equations. This is due to the fact that the convolutional structure of the Fourier representation of the nonlinear term is the same in both cases and that the Euler equations conserve the solenoidal component of the $L^{2}$ norm for smooth initial data.
@@
3. Numerical Methods
In this section we will detail the numerical methods we utilize to simulate Equations 2.6 and 2.7. All simulations depicted in this study were performed Matlab (R2023a) using our own code. For this computational study we implemented the KdV equation on $\mathbf{R}$, subject to periodic B.C. with period 2 with pseudo-spectral methods. Specifically, we implement the equations using an integrating factor method to compute the linear term exactly with a fourth-order Runge-Kutta scheme for the timestepping with the nonlinear term computed explicitly. For an overview of integrating factor schemes see e.g. [57, 77] and the references contained within. Our particular choice of numerical scheme, which is given Algorithm 1, was chosen as it allows for larger timesteps than traditional schemes for KdV, such as the Zabusky-Kruskal explicit leapfrog scheme seen in the literature [79].
We note that in the simulations of Equation 2.6, depicted below, we utilized different spatial resolutions depending on the particular choice of initial profile and dispersion coefficient, $\delta^{2}$. We ensured in our simulations that $\Delta t\lesssim{\Delta x}^{2}$, which was consistent with the implementation of this timestepping scheme in chapter 10 of [77] and appeared to be numerically stable for all spatial resolutions we utilized in this work. Here $\Delta t=t_{n+1}-t_{n}$ and $\Delta x=\frac{2}{N}$ represent the temporal and spatial resolution scales with $N$ indicating the spatial resolution, given in powers of $2$. While we do not explicitly state the spatial resolution for each simulation, one can see in the spectral plots, i.e. Figures 1, 2, 3, 4, 5 and 6, that the x-axis covers $[1,N/2]$. In this study, $N$ ranges between $2^{7}$ and $2^{11}$, depending on the amount of Fourier modes required to accurately simulate a given solution. $N$ is chosen specifically to ensure that the solution is well resolved, which here means that Fourier coefficients of the reference solution decay to machine precision before the dealiasing cutoff, i.e. the vertical red line featured in Figures 1, 2, 3, 4, 5 and 6. That is, the momentum decays exponentially in wave number, we take $N$ large enough such that the momentum decays to machine precision (given by $162.216e-$ 16) before the wave number $\frac{N}{3}$. We utilize 2/3 truncation for dealiasing to accurately calculate the nonlinear term, i.e., we set the last 1/3 of the Fourier modes to zero to prevent aliasing effects. Additionally, we note that while the red vertical line on spectral plots (see Figures 6(b), 6(d), 4(b) and 5(b) denotes the dealiasing cutoff, the vertical blue line denotes the observational window cutoff, that is, the Fourier modes to the left of and including the blue vertical line are accurately observed and used in the computation of the nudging term.
@@ 3. Numerical Methods
1:Observational operator $H:=P_{M}$ (projection onto the lowest $M$ Fourier modes),
@@
4:Linear function $L:=-\delta^{2}ik^{3}+\gamma$ for integrating factor,
5:${u}_{n}$ and ${v}_{n}$, the value of solutions $u$ and $v$ at previous iteration,
6:$\hat{u}_{n}=\mathcal{F}(u_{n})$ and $\hat{v}_{n}=\mathcal{F}(v_{n})$ represent the Fourier transforms of $u_{n}$ and $v_{n}$, respectively. Note that $\hat{v}_{n}(k)$ is a function of the wave index, $k$, and the indices present here indicate the current timestep.
7:Forward time integration $\hat{u}_{n+1}:=M(\hat{u}_{n}).$
@@
The nudging term is implemented numerically in Algorithm 1 as an explicit term. This implementation of the feedback-control term that we used induces a CFL constraint required for the numerical scheme to remain stable, $\mu\lesssim\frac{2}{\Delta t}$. This CFL constraint can be removed by utilizing an implicit formulation of the nudging term, which we do not consider in this work. It is worth noting that the nudging term cannot be implemented numerically using multi-stage methods, such as RK4, due to the need to interpolate in time (see e.g. [68]). The simulations featured in this work primarily use a value of $\mu=100$, which was found to be sufficient for this study. For an in-depth examination of the role the value of the nudging parameter plays in convergence, see [15].
@@
To test the effectiveness of nudging on the KdV equation Equation 2.6 we utilized an “identical twin” experimental design. This experimental design is standard in the validation of data assimilation methods, where we first initialize a reference solution. This solution will be treated as the true solution which we will attempt to recover using observations. The observations will be given by the lowest $M$ Fourier modes. We note that the observations are considered to be perfect with no distortions (i.e., no error measurements) or stochastic noise.
@@
We note that typically in the data assimilation literature for dissipative evolutionary equations, the equations are evolved forward in time long enough such that the solution remains in an absorbing set in a relevant phase space. That is, we run solutions long enough forward in time such that they are uniformly bounded with respect to time in $L^{2}$, $H^{1}$, or another relevant norm for all future times, see e.g., [74, 75] for a full definition of absorbing sets as well as their properties. In this study we do not evolve the equations forward in time to arbitrarily large times, as solutions are not dissipative and so there is not necessarily an absorbing set that attracts all bounded solutions after a long enough time. In fact, as the $L^{2}$ norm is a conserved quantity for Equation 2.6, such an absorbing set cannot exist in $L^{2}$. We opted to simply evolve forward the initial profile a given amount of time, typically $t=1$, in order to ensure that the reference solution has developed interesting dynamics outside the observation domain. While solutions can develop small length scales not captured by the observed modes, they typically have a finite number of active Fourier modes, with the total number of modes active at a given time dependent on the initial condition, the current time, the total momentum in the system, and the dispersion parameter, $\delta^{2}$. An active Fourier mode in this sense means that |$\absolutevalue{\hat{u}_{k}}>\epsilon$, where $k\in\mathbf{Z}$ is the wavemode index and $\epsilon$ indicates machine precision.
4. Computational Results
4.1. Classical KdV equation
@@
In this section we will examine how the data assimilation nudging algorithm fails for Equation 2.6. Nudging appears to fail subtly, with the extent to which it works appears to be determined by whether the active Fourier modes of the initial profile are contained in the annulus of the observed Fourier modes, as we will see later on in Figure 5. We will start by considering examples such as in 2.0.5 to illustrate how nudging fails.
@@
where $c\in\mathbf{R}$ is an arbitrary constant with $c\neq 0$. Note that by 2.0.3 we know that as this solution is initially $\frac{2}{k_{0}}$ periodic, it will remain $\frac{2}{k_{0}}$ periodic for all time $t\in\mathbf{R}$.
@@
Using initial data of the form Equation 4.1 we can now construct solutions to Equation 2.6 which the nudging algorithm Equation 2.7 explicitly fails to recover. Given a number of observed modes $M$, we can set $k_{0}\in\mathbf{Z}$ such that |$\absolutevalue{k_{0}}\geq M+1$. Again, we emphasize that solutions initialized in this way (see Equation 4.1) are initially $\frac{2}{k_{0}}$ periodic and remain $\frac{2}{k_{0}}$ periodic due to 2.0.3. Noting that $M<k_{0}$, it follows immediately from 2.0.5 that $P_{M}u(t)\equiv 0$ where $u$ is the solution of Equation 2.6 generated from the initial data prescribed by Equation 4.1. This shows that a fixed number of observed Fourier modes are insufficient to distinguish the zero solution from solutions initialized with suitable high frequency oscillations, and so one should not expect data assimilation to succeed. Indeed, if we choose $v$ to be the solution of Equation 2.7 initialized by the identically zero solution, then it is easy to deduce that $P_{M}(u-v)\equiv 0$ holds for all time and so the solution to Equation 2.7 simply corresponds to a solution to Equation 2.6 albeit with a different initial condition than the reference solution.
@@
One can see computationally in Figure 1 that nudging fails to recover the solution $u(x,t)$, of Equation 2.6 generated from the initial data prescribed by Equation 4.1. In Figure 1 the solution, $v(x,t)$, of Equation 2.7 is initialized with the identically zero solution. Exactly as one expects, we find that $v\equiv 0$ holds for the entire simulation. Similarly, we see in Figure 2 the roles of $u$ and $v$ are reversed. That is, we instead initialize the solution to $v$ with Equation 4.1 and the reference solution, $u$, is initialized with the identically zero solution. We see in Figure 2 similar behavior occurs, as the feedback-control term remains zero for the entire simulation and both equations evolve as solutions of Equation 2.6 with different initial conditions. While this is expected, it is nevertheless an interesting result as it shows that the initialization of the assimilated solution to Equation 2.7 matters. This fact is fundamentally different to many analytical results regarding the application of the nudging algorithm to dissipative equations, as for dissipative systems the effect of the initial condition fades given enough time, contrary to what occurs in this case.
@@
The previous counterexample illustrates that data assimilation methods cannot uniquely determine the observed reference solution when the observations are identically zero, but one may hope that more physically realistic (i.e. not identically zero) observations will yield better results. This is true to some extent, however the level of success is highly dependent on the initial profile of the reference solution and one can similarly construct solutions that exhibit high frequency oscillations that nudging cannot recover. If given an analytic solution $u$ in $L^{2}_{\text{per}}$ of Equation 2.6, we can indeed construct infinitely many solutions where nudging fails to synchronize with the reference solution given a finite number of observed Fourier modes.

### v2

@@ 3.1. Preliminaries and Analytical Results
Moreover, because of the $\frac{L}{k}$ periodicity of $f(x)$ one has that $v(x,t)$ is also a solution of Equation˜24. As $v(x,0)=u^{in}(x)$, then by the uniqueness of solutions to Equation˜24 we obtain that $v(x,t)=u(x,t)$, hence we have that $u(x+\frac{L}{k},t)=u(x,t)$. The fact that $u$ restricted to [T,T][-T,T] is continuous and bounded follows immediately from Theorem 10 in [12]. ∎
Corollary 3.4.
Under the conditions of Proposition˜3.3 the space of periodic functions with period $\frac{L}{k}$ is invariant under the solution of Equation˜24.
Proposition 3.5.
Suppose $k\in\mathbb{Z}\setminus\{0\}$ is given and suppose $u^{in}_{j}\in H^{2}_{\text{per}}$, for $j=1,2$, satisfying
$u^{in}_{j}(x+\frac{L}{k})=u_{j}^{in}(x)$ (25)
for a.e. $x\in\mathbb{R}$ and $u_{j}^{in}$ has spatial mean zero. Moreover, assume $\mathinner{\!\left\lVert u^{in}_{1}\right\rVert}_{L^{2}}\neq\mathinner{\!\left\lVert u^{in}_{2}\right\rVert}_{L^{2}}$ and let $u_{j}(x,t)$ be the corresponding solutions of Equation˜19. Then
@@
    $u_{j}(x+\frac{L}{k},t)=u_{j}(x,t)$ for a.e. $x\in\mathbb{R}$ and for $j=1,2$ and all $t\in\mathbb{R}$.
@@
    $P_{M}u_{1}(\cdot,t)=P_{M}u_{2}(\cdot,t)$ for any positive integer $M<\mathinner{\!\left\lvert k\right\rvert}$, and every $t\in\mathbb{R}$.
@@
    $\limsup\limits_{t\to\infty}\mathinner{\!\left\lVert u_{1}(\cdot,t)-u_{2}(\cdot,t)\right\rVert}_{L^{2}}>0.$
@@
We note that as Equation˜24 is a particular case of Equation˜24 (i.e. $\gamma=0,\beta=0$), then (i) follows immediately from Proposition˜3.3. Part (ii) then follows immediately from part (i) combined with Proposition˜3.2.
Let us now prove part (iii). Let $u_{1}$ and $u_{2}$ be two solutions to Equation˜19 under the assumptions of the proposition. Without loss of generality, suppose $\mathinner{\!\left\lVert u_{1}^{\text{in}}\right\rVert}_{L^{2}}>\mathinner{\!\left\lVert u_{2}^{\text{in}}\right\rVert}_{L^{2}}$. We note that the $L^{2}$ norm is a conserved quantity for Equation˜19 and thus for $j=1,2$ we have that $\mathinner{\!\left\lVert u_{j}(t)\right\rVert}_{L^{2}}=\mathinner{\!\left\lVert u_{j}^{\text{in}}\right\rVert}_{L^{2}}$ holds for all $t\in\mathbb{R}$. The result follows from an application of the reverse triangle inequality as
$0<\mathinner{\!\left\lVert u_{1}(\cdot,t)\right\rVert}_{L^{2}}-\mathinner{\!\left\lVert u_{2}(\cdot,t)\right\rVert}_{L^{2}}\leq\mathinner{\!\left\lVert u_{1}(\cdot,t)-u_{2}(\cdot,t)\right\rVert}_{L^{2}}$ (26)
holds for all time $t\in\mathbb{R}$.
Remark 3.6.
Proposition˜3.5 can be adjusted without change to allow $u_{1}$ and $u_{2}$ to be initialized with periods $\frac{L}{k_{1}}$ and $\frac{L}{k_{2}}$, respectively. The only changes in this case are that $u_{j}$ remain $\frac{2}{k_{j}}$ periodic for all $t\in\mathbb{R}$ in (i) and $M<\min\{\mathinner{\!\left\lvert k_{1}\right\rvert},\mathinner{\!\left\lvert k_{2}\right\rvert}\}$ is required for (ii).
Now, using these propositions, we show that Equation˜19 does not have the finitely many determining modes property.
Theorem 3.7.
The Korteweg-de Vries equation Equation˜19 subject to periodic boundary conditions with period $L$ does not possess the finitely many determining modes property in $L^{2}$.
@@
The proof of this follows straightforwardly from Proposition˜3.5. Assume, by way of contradiction, that the KdV equation enjoys the finitely many determining Fourier modes property in $L^{2}_{\text{per}}$. That is, there exists some $M\in\mathbb{N}$ such that for any two solutions $u_{1}$ and $u_{2}$ to Equation˜19 that if
$\lim_{t\to\infty}\mathinner{\!\left\lVert P_{M}(u_{1}(t)-u_{2}(t))\right\rVert}_{L^{2}}=0,$ (27)
@@
$\lim_{t\to\infty}\mathinner{\!\left\lVert u_{1}(t)-u_{2}(t)\right\rVert}_{L^{2}}=0.$ (28)
Now, we will simply use $k=M+1$, and consider $u_{1}$ and $u_{2}$ to be any two solutions to Equation˜19 such that $u_{1}$ and $u_{2}$ both are smooth enough and have zero spatial mean, are $\frac{L}{k}$ periodic initially, and $\mathinner{\!\left\lVert u_{1}^{in}\right\rVert}_{L^{2}}\neq\mathinner{\!\left\lVert u_{2}^{in}\right\rVert}_{L^{2}}$. We note that two such solutions must exist, as one could take $u_{1}$ to be the zero solution and $u_{2}$ to be the solution generated from the initial data $u_{2}^{in}=\cos(\frac{2\pi}{L}kx).$
Now, we immediately obtain a contradiction using Proposition˜3.5. ∎
Remark 3.8.
The proof above can be adapted almost without change to the unforced 2D incompressible Euler equations and other non-dissipative systems with a similar nonlinearity. This is due to the fact that the convolutional structure of the Fourier representation of the nonlinear term is the same in both cases and that the Euler equations conserve the solenoidal component of the $L^{2}$ norm for smooth initial data.
@@
3.2. Computational Results
In this section we will detail the numerical methods we utilize to simulate Equations˜19 and 20. All simulations depicted in this study were performed Matlab (R2023a) using our own code. For this computational study we implemented the KdV equation on $\mathbb{R}$, subject to periodic B.C. with period 2 with pseudo-spectral methods. Specifically, we implement the equations using an integrating factor method to compute the linear term exactly with a fourth-order Runge-Kutta scheme for the timestepping with the nonlinear term computed explicitly. For an overview of integrating factor schemes see e.g. [55, 78] and the references contained within. Our particular choice of numerical scheme, which is given Algorithm˜1, was chosen as it allows for larger timesteps than traditional schemes for KdV, such as the Zabusky-Kruskal explicit leapfrog scheme seen in the literature [80].
We note that in the simulations of Equation˜19, depicted below, we utilized different spatial resolutions depending on the particular choice of initial profile and dispersion coefficient, $\delta^{2}$. We ensured in our simulations that $\Delta t\lesssim{\Delta x}^{2}$, which was consistent with the implementation of this timestepping scheme in chapter 10 of [78] and appeared to be numerically stable for all spatial resolutions we utilized in this work. Here $\Delta t=t_{n+1}-t_{n}$ and $\Delta x=\frac{2}{N}$ represent the temporal and spatial resolution scales with $N$ indicating the spatial resolution, given in powers of $2$. While we do not explicitly state the spatial resolution for each simulation, one can see in the spectral plots, i.e. Figures˜4, 5, 6, 7, 8 and 9, that the x-axis covers $[1,N/2]$. In this study, $N$ ranges between $2^{7}$ and $2^{11}$, depending on the amount of Fourier modes required to accurately simulate a given solution. $N$ is chosen specifically to ensure that the solution is well resolved, which here means that Fourier coefficients of the reference solution decay to machine precision before the dealiasing cutoff, i.e. the vertical red line featured in Figures˜4, 5, 6, 7, 8 and 9. That is, the momentum decays exponentially in wave number, we take $N$ large enough such that the momentum decays to machine precision (given by 2.216e162.216e-16) before the wave number $\frac{N}{3}$. We utilize 2/3 truncation for dealiasing to accurately calculate the nonlinear term, i.e., we set the last 1/3 of the Fourier modes to zero to prevent aliasing effects. Additionally, we note that while the red vertical line on spectral plots (see Figures˜9(b), 9(d), 7(b) and 8(b) denotes the dealiasing cutoff, the vertical blue line denotes the observational window cutoff, that is, the Fourier modes to the left of and including the blue vertical line are accurately observed and used in the computation of the nudging term.
@@ 3.2. Computational Results
1:Observational operator $H\mathrel{\mathop{\ordinarycolon}}=P_{M}$ (projection onto the lowest $M$ Fourier modes),
@@
4:Linear function $L\mathrel{\mathop{\ordinarycolon}}=-\delta^{2}ik^{3}+\gamma$ for integrating factor,
[…]
7:Forward time integration $\hat{u}_{n+1}\mathrel{\mathop{\ordinarycolon}}=M(\hat{u}_{n}).$
@@
The nudging term is implemented numerically in Algorithm˜1 as an explicit term. This implementation of the feedback-control term that we used induces a CFL constraint required for the numerical scheme to remain stable, $\mu\lesssim\frac{2}{\Delta t}$. This CFL constraint can be removed by utilizing an implicit formulation of the nudging term, which we do not consider in this work. It is worth noting that the nudging term cannot be implemented numerically using multi-stage methods, such as RK4, due to the need to interpolate in time (see e.g. [67]). The simulations featured in this work primarily use a value of $\mu=100$, which was found to be sufficient for this study. For an in-depth examination of the role the value of the nudging parameter plays in convergence, see [15].
@@
To test the effectiveness of nudging on the KdV equation Equation˜19 we utilized an “identical twin” experimental design. This experimental design is standard in the validation of data assimilation methods, where we first initialize a reference solution. This solution will be treated as the true solution which we will attempt to recover using observations. The observations will be given by the lowest $M$ Fourier modes. We note that the observations are considered to be perfect with no distortions (i.e., no error measurements) or stochastic noise.
@@
We note that typically in the data assimilation literature for dissipative evolutionary equations, the equations are evolved forward in time long enough such that the solution remains in an absorbing set in a relevant phase space. That is, we run solutions long enough forward in time such that they are uniformly bounded with respect to time in $L^{2}$, $H^{1}$, or another relevant norm for all future times, see e.g., [74, 76] for a full definition of absorbing sets as well as their properties. In this study we do not evolve the equations forward in time to arbitrarily large times, as solutions are not dissipative and so there is not necessarily an absorbing set that attracts all bounded solutions after a long enough time. In fact, as the $L^{2}$ norm is a conserved quantity for Equation˜19, such an absorbing set cannot exist in $L^{2}$. We opted to simply evolve forward the initial profile a given amount of time, typically $t=1$, in order to ensure that the reference solution has developed interesting dynamics outside the observation domain. While solutions can develop small length scales not captured by the observed modes, they typically have a finite number of active Fourier modes, with the total number of modes active at a given time dependent on the initial condition, the current time, the total momentum in the system, and the dispersion parameter, $\delta^{2}$. An active Fourier mode in this sense means that $\mathinner{\!\left\lvert\hat{u}_{k}\right\rvert}>\varepsilon$, where $k\in\mathbb{Z}$ is the wavemode index and $\varepsilon$ indicates machine precision.
3.2.1. Classical KdV equation
@@
In this section we will examine how the data assimilation nudging algorithm fails for Equation˜19. Nudging appears to fail subtly, with the extent to which it works appears to be determined by whether the active Fourier modes of the initial profile are contained in the annulus of the observed Fourier modes, as we will see later on in Figure˜8. We will start by considering examples such as in Proposition˜3.5 to illustrate how nudging fails.
@@
where $c\in\mathbb{R}$ is an arbitrary constant with $c\neq 0$. Note that by Proposition˜3.3 we know that as this solution is initially $\frac{2}{k_{0}}$ periodic, it will remain $\frac{2}{k_{0}}$ periodic for all time $t\in\mathbb{R}$.
@@
Using initial data of the form Equation˜29 we can now construct solutions to Equation˜19 which the nudging algorithm Equation˜20 explicitly fails to recover. Given a number of observed modes $M$, we can set $k_{0}\in\mathbb{Z}$ such that $\mathinner{\!\left\lvert k_{0}\right\rvert}\geq M+1$. Again, we emphasize that solutions initialized in this way (see Equation˜29) are initially $\frac{2}{k_{0}}$ periodic and remain $\frac{2}{k_{0}}$ periodic due to Proposition˜3.3. Noting that $M<k_{0}$, it follows immediately from Proposition˜3.5 that $P_{M}u(t)\equiv 0$ where $u$ is the solution of Equation˜19 generated from the initial data prescribed by Equation˜29. This shows that a fixed number of observed Fourier modes are insufficient to distinguish the zero solution from solutions initialized with suitable high frequency oscillations, and so one should not expect data assimilation to succeed. Indeed, if we choose $v$ to be the solution of Equation˜20 initialized by the identically zero solution, then it is easy to deduce that $P_{M}(u-v)\equiv 0$ holds for all time and so the solution to Equation˜20 simply corresponds to a solution to Equation˜19 albeit with a different initial condition than the reference solution.
@@
One can see computationally in Figure˜4 that nudging fails to recover the solution $u(x,t)$, of Equation˜19 generated from the initial data prescribed by Equation˜29. In Figure˜4 the solution, $v(x,t)$, of Equation˜20 is initialized with the identically zero solution. Exactly as one expects, we find that $v\equiv 0$ holds for the entire simulation. Similarly, we see in Figure˜5 the roles of $u$ and $v$ are reversed. That is, we instead initialize the solution to $v$ with Equation˜29 and the reference solution, $u$, is initialized with the identically zero solution. We see in Figure˜5 similar behavior occurs, as the feedback-control term remains zero for the entire simulation and both equations evolve as solutions of Equation˜19 with different initial conditions. While this is expected, it is nevertheless an interesting result as it shows that the initialization of the assimilated solution to Equation˜20 matters. This fact is fundamentally different to many analytical results regarding the application of the nudging algorithm to dissipative equations, as for dissipative systems the effect of the initial condition fades given enough time, contrary to what occurs in this case.
@@
The previous counterexample illustrates that data assimilation methods cannot uniquely determine the observed reference solution when the observations are identically zero, but one may hope that more physically realistic (i.e. not identically zero) observations will yield better results. This is true to some extent, however the level of success is highly dependent on the initial profile of the reference solution and one can similarly construct solutions that exhibit high frequency oscillations that nudging cannot recover. If given an analytic solution $u$ in $L^{2}_{\text{per}}((0,L))$ of Equation˜19, we can indeed construct infinitely many solutions where nudging fails to synchronize with the reference solution given a finite number of observed Fourier modes.